OpenStudy (erinweeks):

Use Gaussian elimination to find the complete solution to the system of equations, or state that none exists.

OpenStudy (erinweeks):

3x - 2y + 2z - w = 2 4x + y + z + 6w = 8 -3x + 2y - 2z + w = 5 5x + 3z - 2w = 1

OpenStudy (zzr0ck3r):

well, this is much more involved and would take a very long time for me to draw it all out. I think it would be easier to see a few examples and a walk-threw

OpenStudy (erinweeks):

A. (2, 0, 3/37, 9/37) B. ( 1, -1/3, 4/9, 6) C. ∅ D. ( 1/2, 0, -37/3, 37/9)

OpenStudy (erinweeks):

Okay if that will help me!

OpenStudy (zzr0ck3r):

you could plug in each solution to check

OpenStudy (zzr0ck3r):

so make x = 2,y=0,....

OpenStudy (zzr0ck3r):

in the long run it is exactly this elimination method for 2 equations with 2 variables y+3x=7 2y+2x=7 we multiply the first equation by -2 -2y-6x=-14 2y+2x=7 now we add the equations -4x=-7 x=7/4

OpenStudy (erinweeks):

i already tried that and i got nowhere

OpenStudy (zzr0ck3r):

its the same thing but with n equations and n variables

OpenStudy (zzr0ck3r):

well if none of the solutions made sense when you pluged them in, then the system has no solutions, and thus we choose c the nullset

OpenStudy (erinweeks):

well i wanted someone to basically make sure of my answer!!!

OpenStudy (erinweeks):

@ganeshie8 helpp please :/

ganeshie8 (ganeshie8):

wolfram is the best thing to double confirm ur work : it says the same - No soultion exist http://www.wolframalpha.com/input/?i=Solve+%3A+3x+-+2y+%2B+2z+-+w%09%3D%092%2C++4x+%2B+y+%2B+z+%2B+6w%09%3D%098%2C++-3x+%2B+2y+-+2z+%2B+w%09%3D%095%2C++5x+%2B+3z+-+2w%09%3D%091

ganeshie8 (ganeshie8):

But it is good to get hang of gaussian elimination as it is useful when u deal wid finidng inverses and stuff..

OpenStudy (erinweeks):

oh gosh what would i do without @ganeshie8 (:

OpenStudy (zzr0ck3r):

for sure it did not want you to just check the solutions it wants you to use elimination. you should take the time to learn this it really is just adding and multiplying easy numbers (lots of them sometimes) but you will need to know how to do it at some point.

ganeshie8 (ganeshie8):

^

OpenStudy (zzr0ck3r):

hmm that link sucks, ignore it and ill find a better one.

OpenStudy (zzr0ck3r):

http://linear.ups.edu/jsmath/latest/fcla-jsmath-latestli18.html this is good. I think you should read it all. you can pm me if you have questions.